OpenStudy (anonymous):

Factor completely 3x2 + 9x − 54. (3x − 9)(x + 6) 3(x − 3)(x + 6) (3x − 6)(x + 9) 3(x − 2)(x + 9)

Directrix (directrix):

For the first step, factor out the common factor of 3. 3^x2 + 9x − 54 = 3* (x^2 + 3x - 18)

Directrix (directrix):

Focus on factoring: (x^2 + 3x - 18) for now.

Directrix (directrix):

What are two numbers that multiply to -18 and also add to +3. We need those to help factor. @aep0814

OpenStudy (anonymous):

-3 and 6

Directrix (directrix):

Yes. x^2 + 3x - 18 = x^2 + 6x -3x -18 The +3x has been rewritten as 6x - 3x

Directrix (directrix):

Factoring by Grouping: x^2 + 6x -3x -18 = x*(x + 6) - 3* (x + 6) = Note the common factor of (x + 6) Factor it out (x + 6) * (x - 3) But, that's not the answer.

Directrix (directrix):

Remember the 3 we factored out at the beginning: 3^x2 + 9x − 54 = 3* (x^2 + 3x - 18) = 3* (x + 6) * (x - 3) is the complete factorization. @aep0814

Directrix (directrix):

Question about this?

OpenStudy (anonymous):

No, thank you so much! :)